# Euclid: Early Release Observations – The star cluster systems of the Local Group dwarf galaxies IC 10 and NGC 6822††thanks: This paper is published on behalf of the Euclid Consortium

### v1

@@ 5.1 Cluster photometry
To go from aperture magnitudes to total magnitudes, aperture corrections are required. For point sources, the correction can be determined in a straightforward manner from the instrumental PSF, while for extended sources, we must also consider the intrinsic radial profile. Unfortunately, this represents a major source of uncertainty in star cluster studies (e.g., Chandar et al. 2010) and methods to deal with this have been widely discussed in the literature (e.g., Cook et al. 2019; Adamo et al. 2015; Deger et al. 2022). A common approach is to derive a relationship between the slope of the enclosed flux growth curve at a given radius and the aperture correction, which can be calibrated using either isolated real clusters or injected synthetic ones. We again followed the approach of L15, who sampled the slope of the growth curve at radii where it is typically at its steepest, and calibrated the relationship using a small sample of relatively bright and isolated star clusters in each galaxy (clusters ESCC-IC10-01, -03, -17, -45, -49, -50, and -52, and ESCC-NGC6822-05, -07, -08, -15, and -28).
@@
Tables 9 and 9 present the photometric measurements for all candidates with classes 1–4 and will be made available electronically. In NGC 6822, the resulting median photometric uncertainties in the V- and -bands are 0.02 and 0.04 for the brightest 50% of clusters, increasing to 0.08 and 0.13 among the faintest half. In IC 10, the corresponding photometric uncertainties are 0.05 and 0.06 for the brighter clusters, rising to 0.08 and 0.22 in the fainter population. Some clusters were particularly faint in some photometric bands or partially obscured by nearby bright stars, resulting in large photometric errors. These clusters are excluded from our analysis and flagged in our photometry tables with an asterisk.
@@ 5.2 SED fitting
To derive the physical properties (e.g., ages, masses, metallicities) as well as the line-of-sight extinction values of the clusters, we compare our photometric UBVRI + measurements to spectral energy distributions (SEDs) predicted by the 2016 version of Bruzual & Charlot (2003) models, which are updated using the stellar evolutionary tracks of Bressan et al. (2012) and Marigo et al. (2013). We do not include in SED fitting, as it offers no additional information compared to the optical bands. The stellar population models are constructed and fit using the BAGPIPES code (Carnall et al. 2018). Prior to fitting, Euclid AB magnitudes and UBVRI Vega magnitudes were converted to flux using the zero points provided by the SVO filter profile service. The BAGPIPES code assumes a fully sampled initial mass function from Kroupa (2001) and we assume a single burst of star formation can approximate all clusters. BAGPIPES uses Bayesian inference to derive the properties of the clusters, constructing a probability distribution based upon the value for each model fit (the likelihood), assuming uncertainties are Gaussian and independent. Priors for each parameter can also be set, which we discuss below and detail in Table 5. The posterior distribution is sampled using the MULTINEST nested sampling algorithm (Skilling 2006; Feroz & Hobson 2008; Feroz et al. 2009) via the PYMULTINEST interface (Buchner et al. 2014). The median of the posterior distribution (50th percentile) is used to determine the best estimate of each cluster property, with the 16th and 84th percentiles defining the uncertainties. To deal with dust, we adopt a Cardelli et al. (1989) extinction law, an value of 3.1 and allow the total line-of-sight extinction in the V-band, , to vary according to the range of values in the literature (detailed in Sects. 2.1 and 2.2). We do not correct for foreground extinction before fitting, as BAGPIPES assumes a foreground screen approximation, and applying a single extinction law consistently avoids introducing additional assumptions about the foreground and the intrinsic source spectrum used to compute values. Although the internal extinction law in dwarf galaxies may differ from that of the MW, the uncertainties involved in estimating foreground extinction and correcting for it separately are significant enough that we opt to model the total extinction with a single component. Within a central 3 radius of IC 10, we impose an prior with a minimum of 0.8, reflecting the higher extinction values reported in the literature. Beyond this radius, the prior is allowed to extend down to 0.65. For NGC 6822, the prior ranges from 0.2 to 0.45 in the central regions (approximately within a radius of 65), and is capped at 0.35 beyond this, consistent with previous studies. Additionally, we applied high-metallicity, low-age priors to five clusters in NGC 6822 and seven in IC 10 that appeared as prominent sources in continuum-subtracted images that we constructed from the LGGS images. Specifically, for these clusters we limit the fits to metallicity values of 0.2–0.3 for NGC 6822 and 0.2–0.5 for IC 10, and ages of 0–10 Myr. We note that since most of the clusters to be fit did not have strong coincident emission, we did not include a nebular component in the fits. Tables 11 and 11 present the BAGPIPES SED-fitting results for all candidates with classes 1–4 and will be made available electronically.
@@ 6 Validation and completeness
Clusters were generated across a grid of ages (10 Myr to 10 Gyr) and total stellar masses ( to ). Metallicity was fixed at , and a Kroupa initial mass function (IMF) was assumed (Kroupa 2001).
@@
Example synthetic clusters, spanning a range of physical properties, are demonstrated in Fig. 4. To make the analysis more tractable, we focused on inserting clusters into a representative inner and an outer region of each galaxy, selected from the full Euclid FoV. The inner regions cover the dense central areas of the galaxies, defined approximately by a box of width 15 centred on NGC 6822 and 12 in IC 10. The outer regions, located beyond these boxes, probe the sparser halo environments. In all, clusters were inserted at random positions in each region (around 1,400 clusters in total). We then performed a blind search for the synthetic clusters and recorded which ones were successfully recovered.
The left panel of Fig. 5 shows the completeness fraction, defined as the recovery rate, versus the magnitude, for each region and galaxy. It shows the expected behaviour of high completeness at bright magnitudes, falling to lower values for fainter clusters. We estimate the and completeness limits by fitting a logistic function to the decay using a non-linear least squares method; the resulting limits are listed in Table 6. These magnitude limits are assumed to be constant with age. When translating these magnitude limits into mass limits (see Sect. 7.2), we find that, overall, we are complete to for ages , and to for ages of , in both galaxies. We note, however, that the assumption of constant magnitude completeness limits is a simplification, as completeness may vary somewhat with age due to differences in cluster appearance.
@@
| NGC 6822 inner | 19.7 | 18.4 |
| NGC 6822 outer | 21.7 | 20.1 |
| IC 10 inner | 21.1 | 19.0 |
| IC 10 outer | 22.1 | 20.3 |
We also examine how the completeness depends on half-light radius, as shown in the right-hand panels of Fig. 5. Our results indicate that the completeness increases with half-light radius in both the inner and outer regions of the galaxies, reaching a maximum at cluster half-light radii of 2 and 4 pc in the inner regions. Beyond this range, recovery rates in the inner regions begin to decline, while in the outer regions, they continue to rise. In the inner regions, very few clusters are recovered with half-light radii exceeding 8 pc due to the difficulty discerning them against the ambient stellar field.
@@
| ID | This study | Veljanoski et al. (2015) | Hwang et al. (2011) | Huxor et al. (2013) | Krienke & Hodge (2004) |
@@
A detailed comparison of our derived masses and ages with those from L15 is provided in Fig. 9. Both studies utilise the Bruzual & Charlot (2003) models but differences stem from the version used, variations in SED-fitting techniques, prior selections, and the inclusion of NIR bands in our analysis. Although we observe fairly good agreement in mass, discrepancies are noted in age estimates. Notably, a number of clusters for which L15 finds old ages are found to be considerably younger in our work.
@@ 7 Results and discussion
In this section, we present the results of our photometric measurements alongside our SED-fitting results. Unless otherwise specified, we restrict our analyses to clusters with class 1–4 and with photometric errors less than one magnitude in all bands. Our catalogues and results should facilitate a variety of future in-depth studies of the cluster populations in NGC 6822 and IC 10 and we focus only on a few specific science results in this work.
@@ 7.2 Ages and masses of the star clusters
Based on these classifications, we identify eight and 13 young clusters (classes 1–4) in NGC 6822 and IC 10, respectively, along with 28 and 32 intermediate-age clusters, and 16 and 27 old clusters.
@@
Figure 11 shows the cluster mass versus age for both galaxies, along with the 50% completeness limit calculated for the inner regions of each galaxy. This has been derived by converting the magnitude thresholds in Table 6 into corresponding age and mass limits using the models from Bruzual & Charlot (2003). For this, we assume a representative metallicity of along with reddening values of for IC 10 (Massey et al. 2007) and for NGC 6822 (Fusco et al. 2012). Figure 11 also presents masses and ages for a sample of 134 SMC777As a dIrr with a stellar mass of (Pace 2024), the SMC provides a good analog to the systems we study in this paper, although it is a few times more massive. clusters, taken from the study by Gatto et al. (2021). These clusters, primarily situated within the central regions of the SMC (approximately from the centre), represent only a fraction of the cluster candidates in the SMC and Magellanic Bridge compiled by Bica et al. (2020). Gatto et al. (2021) use data from the STEP survey to uniformly derive ages and masses for this sample via colour-magnitude diagram (CMD) fitting, and this homogeneous nature makes it ideal for comparison. That said, it is important to keep in mind that the Gatto et al. (2021) sample is biased in terms of its central location (for example, it does not include NGC 121, which is the only old GC in the SMC). Furthermore, as it is based on clusters that were uncovered in a variety of earlier studies, it is likely to have varying (unquantified) completeness. Nonetheless, due to the close proximity of the SMC, it is expected that it will reach to lower masses at any given age than our Euclid search, as indeed can be seen in Fig. 11.
@@
Both NGC 6822 and IC 10 host clusters that range in mass from to , and with ages from Myr to Gyr, generally consistent with the age range inferred from Fig. 10. We find that older clusters are typically more massive, which probably reflects the preferential disruption and fading of low-mass clusters over time (e.g., Fall & Zhang 2001), as well as a possible preference for the formation of more massive star clusters at early times (e.g., Elmegreen & Efremov 1997; Lahén et al. 2025). On the other hand, clusters younger than a few tens of Myr typically have masses below in both systems. An exception is found in IC 10, which hosts two young, massive clusters: ESCC-IC10-40 (class 1) and ESCC-IC10-71 (class 4), with and , and ages of and Myr, respectively.888L15 discuss a potential Super Star Cluster (SSC) candidate in IC 10, first identified by Hunter (2001) but classified in that study as an OB association. L15 argue that, based on the stellar density in the HST image, this object could be an SSC. We identified this object in our cluster search (ESCC-IC10-44); however, after classification by our group, it was assigned class 5 so very unlikely to be a cluster, consistent with its original designation. Observational evidence suggests that the young cluster mass function follows a power law, with a truncation at the high-mass end that appears to correlate with the SFR surface density () of the host galaxy (Larsen 2009; Adamo et al. 2015; Johnson et al. 2017). Converting the data listed in Table 1 to SFR surface densities within one half-light radius, we find values of and for NGC 6822 and IC 10, respectively. While we do not have enough young clusters in our sample to construct meaningful mass functions, we note that the presence of more massive young star clusters in IC 10 is consistent with this expected dependence of the truncation mass on .
@@
The oldest clusters in the dwarfs reach masses well above , with one exceptional GC in NGC 6822 (SC7, ESCC-NGC6822-08) above even . With a derived mass of , this cluster is roughly twice as massive as any of the other old clusters in NGC 6822 and it is only slightly less massive than Centauri (, van de Ven et al. 2006), the most massive GC in the MW. However, we note that its absolute magnitude () is roughly one magnitude fainter than some estimates of Centauri (; Baumgardt et al. 2020). Interestingly, like Centauri, SC7 is also highly elliptical (see Fig. 1), as originally pointed out by Huxor et al. (2013). Our derived metallicity of is considerably lower than the one found by Larsen et al. 2022 (see Table 7), but in agreement with the metallicity from fitting a deep HST CMD (McGill et al. 2025); Larsen et al. (2022) also find that the cluster has an -abundance close to solar values. The existence of this high ellipticity old massive star cluster in the outskirts of NGC 6822 is particularly intriguing, and it will be examined in further detail in a future paper. While theories for the origin of Centauri often focus on it being the stripped core of a disrupted dwarf galaxy (e.g., Bekki & Freeman 2003), a similar accretion scenario for SC7 may appear less obvious, as NGC 6822 shows no signs of recent tidal disturbance in its outskirts (Zhang et al. 2021). However, we note that the cluster may originate from a much earlier event, now fully phase-mixed and no longer evident in the galaxys present-day structure.

### v3

@@ 5.1 Cluster photometry
To go from aperture magnitudes to total magnitudes, aperture corrections are required. For point sources, the correction can be determined in a straightforward manner from the instrumental PSF, while for extended sources, we must also consider the intrinsic radial profile. Unfortunately, this represents a major source of uncertainty in star cluster studies (e.g. Chandar et al. 2010) and methods to deal with this have been widely discussed in the literature (e.g. Cook et al. 2019; Adamo et al. 2015; Deger et al. 2022). A common approach is to derive a relationship between the slope of the enclosed flux growth curve at a given radius and the aperture correction, which can be calibrated using either isolated real clusters or injected synthetic ones. We again followed the approach of L15, who sampled the slope of the growth curve at radii where it is typically at its steepest, and calibrated the relationship using a small sample of relatively bright and isolated star clusters in each galaxy (clusters ESCC-IC10-01, -03, -17, -45, -49, -50, and -52, and ESCC-NGC6822-05, -07, -08, -15, and -28).
@@
Tables 9 and 9 present the photometric measurements for all candidates with classes 1–4 and are available at the CDS. In NGC 6822, the resulting median photometric uncertainties in the V- and -bands are 0.02 and 0.04 for the brightest 50% of clusters, increasing to 0.08 and 0.13 among the faintest half. In IC 10, the corresponding photometric uncertainties are 0.05 and 0.06 for the brighter clusters, rising to 0.08 and 0.22 in the fainter population. Some clusters were particularly faint in some photometric bands or partially obscured by nearby bright stars, resulting in large photometric errors. These clusters are excluded from our analysis and flagged in our photometry tables with an asterisk.
@@ 5.2 SED fitting
To derive the physical properties (e.g. ages, masses, metallicities) as well as the line-of-sight extinction values of the clusters, we compared our photometric UBVRI + measurements to spectral energy distributions (SEDs) predicted by the 2016 version of Bruzual and Charlot (2003) models, which are updated using the stellar evolutionary tracks of Bressan et al. (2012) and Marigo et al. (2013). We did not include in SED fitting, as it offers no additional information compared to the optical bands. The stellar population models were constructed and fit using the BAGPIPES code (Carnall et al. 2018). Prior to fitting, Euclid AB magnitudes and UBVRI Vega magnitudes were converted to flux using the zero points provided by the SVO filter profile service. The BAGPIPES code assumes a fully sampled initial mass function from Kroupa (2001) and we assumed a single burst of star formation can approximate all clusters. BAGPIPES uses Bayesian inference to derive the properties of the clusters, constructing a probability distribution based upon the value for each model fit (the likelihood), assuming uncertainties are Gaussian and independent. Priors for each parameter can also be set, which we discuss below and detail in Table 5. The posterior distribution was sampled using the MULTINEST nested sampling algorithm (Skilling 2006; Feroz and Hobson 2008; Feroz et al. 2009) via the PYMULTINEST interface (Buchner et al. 2014). The median of the posterior distribution (50th percentile) was used to determine the best estimate of each cluster property, with the 16th and 84th percentiles defining the uncertainties. To deal with dust, we adopted a Cardelli et al. (1989) extinction law, an value of 3.1 and allowed the total line-of-sight extinction in the V-band, , to vary according to the range of values in the literature (detailed in Sects. 2.1 and 2.2). We did not correct for foreground extinction before fitting, as BAGPIPES assumes a foreground screen approximation, and applying a single extinction law consistently avoids introducing additional assumptions about the foreground and the intrinsic source spectrum used to compute values. Although the internal extinction law in dwarf galaxies may differ from that of the MW, the uncertainties involved in estimating foreground extinction and correcting for it separately are significant enough that we opt to model the total extinction with a single component. Within a central 3 radius of IC 10, we imposed an prior with a minimum of 0.8, reflecting the higher extinction values reported in the literature. Beyond this radius, the prior was allowed to extend down to 0.65. For NGC 6822, the prior ranged from 0.2 to 0.45 in the central regions (approximately within a radius of 65), and was capped at 0.35 beyond this, consistent with previous studies. Additionally, we applied high-metallicity, low-age priors to five clusters in NGC 6822 and seven in IC 10 that appeared as prominent sources in continuum-subtracted images that we constructed from the LGGS images. Specifically, for these clusters we limited the fits to metallicity values of 0.2–0.3 for NGC 6822 and 0.2–0.5 for IC 10, and ages of 0–10 Myr. We note that since most of the clusters to be fit did not have strong coincident emission, we did not include a nebular component in the fits. Tables 11 and 11 present the BAGPIPES SED-fitting results (including ages, initial mass formed, living stellar masses, and ) for all candidates with classes 1–4 and are available at the CDS. While BAGPIPES provides both the mass formed and the living stellar mass, it does not have an option for including remnants. However, using FSPS models (Conroy et al. 2009; Conroy and Gunn 2010), we approximate that the inclusion of the latter contributes an additional –5% in present-day mass.
@@ 6 Validation and completeness
Clusters were generated across a grid of ages (10 Myr to 10 Gyr) and stellar masses ( to ). Metallicity was fixed at , and a Kroupa initial mass function (IMF) was assumed (Kroupa 2001).
@@
Example synthetic clusters, spanning a range of physical properties, are demonstrated in Fig. 4. To make the analysis more tractable, we focused on inserting clusters into a representative inner and an outer region of each galaxy, selected from the full Euclid FoV. The inner regions cover the dense central areas of the galaxies, defined approximately by a box of width 15 centred on NGC 6822 and 12 in IC 10. The outer regions, located beyond these boxes, probe the sparser halo environments. In all, clusters were inserted at random positions in each region (around 6000 clusters in total). We then performed a search for the synthetic clusters and recorded which ones were successfully recovered.
Figure 5 shows the completeness fraction, defined as the recovery rate, versus the magnitude and half-light radius for each region and galaxy. The full two-dimensional map is shown alongside the projected completeness fractions shown as histograms along each axis. The projected histograms in the top panels display the expected behaviour of high completeness at bright magnitudes, falling to lower values for fainter clusters. We estimate the and completeness limits by fitting a logistic function to the decay using a non-linear least squares method; the resulting limits are listed in Table 6. These magnitude limits are assumed to be constant with age. When translating these magnitude limits into mass limits (see Sect. 7.2), we find that, overall, we are complete to for ages , and to for ages of , in both galaxies. We note, however, that the assumption of constant magnitude completeness limits is a simplification, as completeness may vary somewhat with age due to differences in cluster appearance.
@@
| NGC 6822 inner | 19.9 | 18.2 |
| NGC 6822 outer | 21.4 | 19.8 |
| IC 10 inner | 20.9 | 18.8 |
| IC 10 outer | 21.9 | 19.7 |
We also examine how the completeness depends on half-light radius, as shown in the vertically-oriented projected histograms in Fig. 5. Our results indicate that the completeness generally decreases with half-light radius across both regions and galaxies, following a maximum at half-light radii of –3 pc. Recovery rates in the inner regions decline steeply beyond this maximum, whereas those in the outer regions decline more shallowly.
@@
| ID | This study | Veljanoski et al. (2015) | Hwang et al. (2011) | Huxor et al. (2013) | Krienke and Hodge (2004) |
@@
A detailed comparison of our derived masses and ages with those from L15 is provided in Fig. 9. Both studies utilise the Bruzual and Charlot (2003) models but differences stem from the version used, variations in SED-fitting techniques, prior selections, and the inclusion of NIR bands in our analysis. Although we observe fairly good agreement in mass, discrepancies are noted in age estimates. Notably, a number of clusters for which L15 finds old ages are found to be considerably younger in our work. We checked if removing the most dust-extincted sources would eliminate this discrepancy, but they persisted, although the overall scatter slightly decreased.
@@ 7 Results and discussion
In this section, we present the results of our photometric measurements alongside our SED-fitting results, tabulated in Appendix B. Unless otherwise specified, we restrict our analyses to clusters with class 1–4 and with photometric errors less than one magnitude in all bands. Our catalogues and results should facilitate a variety of future in-depth studies of the cluster populations in NGC 6822 and IC 10 and we focus only on a few specific science results in this work.
@@ 7.2 Ages and masses of the star clusters
Based on these classifications, we identify eight and 12 young clusters (classes 1–4) in NGC 6822 and IC 10, respectively. We also identify 28 and 32 intermediate-age clusters, and 16 and 27 old clusters in these galaxies.
@@
Figure 11 shows the cluster mass (initial mass formed) versus age for both galaxies, along with the 50% completeness limit calculated for the inner regions of each galaxy. This has been derived by converting the magnitude thresholds in Table 6 into corresponding age and mass limits using the models from Bruzual and Charlot (2003). For this, we assume a representative metallicity of along with reddening values of for IC 10 (Massey et al. 2007) and for NGC 6822 (Fusco et al. 2012). Figure 11 also presents masses and ages for a sample of 134 SMC777As a dIrr with a stellar mass of (Pace 2024), the SMC provides a good analog to the systems we study in this paper, although it is a few times more massive. clusters, taken from the study by Gatto et al. (2021). These clusters, primarily situated within the central regions of the SMC (approximately from the centre), represent only a fraction of the cluster candidates in the SMC and Magellanic Bridge compiled by Bica et al. (2020). Gatto et al. (2021) use data from the STEP survey to uniformly derive ages and masses for this sample via colour-magnitude diagram (CMD) fitting, and this homogeneous nature makes it ideal for comparison. That said, it is important to keep in mind that the Gatto et al. (2021) sample is biased in terms of its central location (for example, it does not include NGC 121, which is the only old GC in the SMC). Furthermore, as it is based on clusters that were uncovered in a variety of earlier studies, it is likely to have varying (unquantified) completeness. Nonetheless, due to the close proximity of the SMC, it is expected that it will reach to lower masses at any given age than our Euclid search, as indeed can be seen in Fig. 11. Both BAGPIPES and the ASteCA (Perren et al. 2015) version used by Gatto et al. (2021) estimate the initial mass formed. While there may be systematic differences between the initial mass computations, these are expected to be small.
@@
Both NGC 6822 and IC 10 host clusters that range in mass from to , and with ages from Myr to Gyr, generally consistent with the age range inferred from Fig. 10. We find that older clusters are typically more massive, which probably reflects the preferential disruption and fading of low-mass clusters over time (e.g. Fall and Zhang 2001), as well as a possible preference for the formation of more massive star clusters at early times (e.g. Elmegreen and Efremov 1997; Lahén et al. 2025). On the other hand, clusters younger than a few tens of Myr typically have masses below in both systems. An exception is found in IC 10, which hosts two young, massive clusters: ESCC-IC10-40 (class 1) and ESCC-IC10-71 (class 4), with and , and ages of and Myr, respectively.888L15 discuss a potential Super Star Cluster (SSC) candidate in IC 10, first identified by Hunter (2001) but classified in that study as an OB association. L15 argue that, based on the stellar density in the HST image, this object could be an SSC. We identified this object in our cluster search (ESCC-IC10-44); however, after classification by our group, it was assigned class 5 so very unlikely to be a cluster, consistent with its original designation. Observational evidence suggests that the young cluster mass function follows a power law, with a truncation at the high-mass end that appears to correlate with the SFR surface density () of the host galaxy (Larsen 2009; Adamo et al. 2015; Johnson et al. 2017). Converting the data listed in Table 1 to SFR surface densities within one half-light radius, we find values of and for NGC 6822 and IC 10, respectively. While we do not have enough young clusters in our sample to construct meaningful mass functions, we note that the presence of more massive young star clusters in IC 10 is consistent with this expected dependence of the truncation mass on .
@@
The oldest clusters in the dwarfs reach masses well above , with one exceptional GC in NGC 6822 (SC7=ESCC-NGC6822-08) above even . With a derived initial mass formed of , this cluster is roughly twice as massive as any of the other old clusters in NGC 6822. Converting this mass to an estimate of the present-day mass (including both living stars and remnants) using FSPS models (Conroy et al. 2009; Conroy and Gunn 2010), we obtain a current mass of . This high mass is further supported by the cluster velocity dispersion reported by Larsen et al. (2022). Using their measurement of 9.9 kms-1, we estimate the dynamical mass to be which is comparable, albeit slightly lower, to our estimated current mass. Our derived metallicity of is considerably lower than the one found by Larsen et al. 2022 (see Table 7), but in agreement with the metallicity from fitting a deep HST CMD (McGill et al., in prep). Aside from its high mass, other peculiarities of SC7 include an -abundance close to the solar value (Larsen et al. 2022) and a high ellipticity (see Fig. 1, and Huxor et al. 2013). The existence of this high ellipticity old massive star cluster in the outskirts of NGC 6822 is particularly intriguing, and it is tempting to speculate that it has been accreted. Indeed, SC7’s high mass and ellipticity are reminiscent of Centauri, the most massive GC in the MW (; van de Ven et al. 2006), which is suspected of being the stripped core of a disrupted dwarf galaxy (e.g. Bekki and Freeman 2003). A similar origin for SC7 may appear less compelling, as NGC 6822 shows no signs of recent tidal disturbance in its outskirts (Zhang et al. 2021). However, we note that the cluster may originate from a much earlier event, now fully phase-mixed and no longer evident in the galaxys present-day structure.